### v1

# The nuclear-spin dark state in silicon

Abstract
We report the formation of a nuclear-spin dark state in a gate-defined silicon double quantum dot. The dark state has been predicted to emerge during dynamic nuclear polarization when the electrons in the quantum dot drive the nuclei in the semiconductor into a decoupled state. We show that, as expected, the electron-nuclear coupling rapidly diminishes in the dark state, and that this state depends on the synchronized precession of the nuclear spins. Moreover, the dark state significantly increases the lifetimes of electronic spin states. This nuclear-spin dark state could be used as a quantum memory or as a resource for quantum sensing, and further studies of the dark state may point the way toward increased polarization of nuclear spin ensembles.
pacs:
I Introduction
Nuclear spins are abundant in condensed matter systems, and the ability to manipulate large ensembles of nuclear spins has led to many technological advances. Nuclear magnetic resonance, for example, has profoundly shaped science and medicine. In the quantum realm, nuclear spin ensembles often create decoherence for semiconductor qubits based on electron spins, and the ability to precisely control these nuclei would be beneficial for applications like quantum computing. Indeed, many ideas from nuclear magnetic resonance have been employed to mitigate this decoherence. One such approach, dynamic nuclear polarization (DNP), aims to suppress fluctuations in nuclear spin ensembles by fully polarizing the nuclei. Platforms including bulk solids [1, 2], gate-defined quantum dots [3, 4, 5], and self-assembled quantum dots [6, 7] all offer routes to DNP, but achieving full polarization remains a challenge. Significant theoretical research has focused on understanding why this is the case. Multiple lines of research have predicted that a collective nuclear-spin dark state, which does not interact with electron spins, should emerge during DNP  [8, 9, 10, 11, 12, 13, 14, 15]. Understanding the nuclear-spin dark state is crucial to improving DNP, and the dark state is also predicted to be an excellent quantum memory and resource for quantum sensing [9, 16, 17].
Despite the importance of the nuclear-spin dark state, only indirect evidence for its existence exists. Data from reconstructed spin states in self-assembled quantum dots are consistent with predicted dark-state properties [18], and measurements of hole-spin polarization recovery in lead-halide perovskite materials have suggested the presence of the dark state [19]. A direct confirmation of the nuclear-spin dark state and tests of its properties would serve as important verifications for these predictions and point the way to harnessing this exotic state for various applications.
Here, we present evidence for the formation of the nuclear-spin dark state in a Si gate-defined quantum dot. We directly probe the electron-nuclear coupling through Landau-Zener sweeps, and we find that it sharply reduces during DNP, in agreement with theoretical predictions. We also observe that the dark state persists for about 1 ms, in agreement with nuclear spin inhomogeneous dephasing times in Si. We also verify that the dark state significantly increases the lifetime of the electronic spin states involved. Together with these experiments, we construct numerical simulations that model the quantum-mechanical behavior of the electrons, as well as the semi-classical behavior of each of the roughly $10^{4}$ nuclear spins interacting with the electrons in our device. Our work verifies key theoretical predictions about the nuclear-spin dark state. Moreover, our work demonstrates that exotic nuclear spin states can occur in the very same devices that are used for quantum computing [20, 21, 22], potentially opening the door to using such states as quantum memories or as resources for quantum sensing or suppressing decoherence.
II Setup
[…]
In this work, we use a gate-defined double quantum dot fabricated on a natural Si/SiGe heterostructure [Fig. 1(a)]. The electron occupancies of dot 1 and dot 2 are controlled by the positive voltages applied to the plunger gates, $V_{p1}$ and $V_{p2}$, respectively. We use the quantum dot beneath gate $P_{cs}$ as a sensor and measure its conductance via rf reflectometry [23] to read out the electronic state in the double dot. The device is cooled in a dilution refrigerator to a temperature of about 10 mK.
We operate the device near the $(4,0)$$(3,1)$ charge transition, with 4(0) or 3(1) electrons in dot 1(2). In this configuration, two electrons in dot 1 always fill the lowest valley level as a singlet. The system dynamics can thus be described within a two-electron picture for the remaining “valence” electrons, which can form an effective singlet-triplet qubit [24, 25]. Figure 1(b) shows the relevant energies of spin states in the $(3,1)$ charge configuration as a function of the detuning $\epsilon$. At a specific detuning $\epsilon_{ST}$, the exchange coupling $J$ equals the electronic Zeeman energy $E_{Z}=\bar{g}\mu_{B}B^{z}_{tot}$, where $\bar{g}\approx 2$ is the electron $g$-factor in Si, and $B^{z}_{tot}$ is the total magnetic field experienced by the electrons, which includes both the externally applied field $B^{z}$ and the longitudinal hyperfine field. At $\epsilon_{ST}$, the spin singlet state, $|S\rangle=(\mid\uparrow\downarrow\rangle-\mid\downarrow\uparrow\rangle)/\sqrt{2}$, comes into resonance with the polarized triplet $|T_{-}\rangle=\mid\downarrow\downarrow\rangle$. The hyperfine interaction between the electron and nuclear spins couples these states. In particular, a difference in the transverse hyperfine field between the two dots couples $S$ and $T_{-}$, and the absolute value of this coupling is denoted by $\Delta_{ST}$. Deep in the (3,1) charge configuration, a difference in the longitudinal hyperfine field couples $S$ and $T_{0}$ [26].
All measurements presented in this paper begin with the state initialized as a singlet in $(4,0)$. Then, we manipulate the spin state of the valence electrons by pulsing $\epsilon$, typically near $\epsilon_{ST}$. To measure the device, we perform a Pauli spin blockade (PSB) readout via radiofrequency reflectometry [23] at $\epsilon=0$ mV, allowing us to distinguish between the singlet and triplet configurations.
Figure 1: Experimental setup. (a) Scanning electron micrograph of a device that is nominally identical to the one used. A double quantum dot is formed under plunger gates $P_{1}$ and $P_{2}$. A sensor dot is formed under gate $P_{cs}$. The magnetic field $B^{z}$ is applied parallel to the axis connecting the two dots, as indicated. (b) Energy level diagram showing the relevant energy levels in the $(3,1)$ charge configuration. The variable $\epsilon$ is defined as a change in the plunger gate voltages such that $(\Delta V_{p1},\Delta V_{p2})=(-\epsilon,\epsilon)$, with $\epsilon=0$ in the PSB region and $\epsilon=1$ mV near the $(4,0)$-$(3,1)$ charge transition. Inset: an LZ sweep through $\epsilon_{ST}$, where the solid (dashed) arrow illustrates a passage through the avoided crossing that does (not) lead to a $S$ to $T_{-}$ transition. (c) Left panel: measurement of the $S\text{-}T_{-}$ avoided crossing position $\epsilon_{ST}$ over time. $\Delta S$ is the measured charge sensor signal relative to the mean signal value at $\epsilon<6.5$ mV. Negative values indicate non-zero triplet probability. Right panel: signal amplitude $|\Delta S|$ at $\epsilon_{ST}$. The data are extracted from the left panel, corresponding to the absolute value of the $\Delta S$ minima for each time value.
III Results
We first demonstrate that DNP occurs in our device. In double quantum dots, DNP can be achieved by bringing $S$ and $T_{-}$ into resonance [4, 5]. A transition from $S$ to $T_{-}$ decreases the $z$-axis angular momentum of the electron spins by $\hbar$. In the absence of effects like spin-orbit coupling, angular momentum in the electron-nuclear system is conserved, and the $z$-axis angular momentum of the nuclear-spin system changes by $\hbar$.
Figure 1(c) shows the results of an experiment with the external magnetic field $B^{z}=1$ mT. We repeatedly initialize the electrons as a singlet and then pulse $\epsilon$ to different values near $\epsilon_{ST}$ for a fixed time of $4~{}\mu$s before the measurement. In the left panel, the thin line indicates the points at which the $S\text{-}T_{-}$ transition occurs [27]. We observe that $\epsilon_{ST}$ increases over time, corresponding to a reduction in $B^{z}_{tot}$. This change in the total field is a hallmark of DNP in double quantum dots and has been observed previously in GaAs double quantum dots [4, 5]. As described in the Supplemental Material, we also observe that the longitudinal magnetic Zeeman gradient $\Delta E_{Z}$ changes during a DNP process, consistent with results from GaAs devices. We do not observe significant DNP effects at larger magnetic fields. This likely occurs because spin-orbit effects [28, 27], which can quench DNP, become significant at higher fields.
Figure 2: Driving into the dark state. (a) Pulse sequence for the experiment. Each pump cycle consists of a singlet load, an LZ sweep through $\epsilon_{ST}$, and a single-shot singlet-triplet measurement. A pause sequence is added after every 64 pump cycles, which pauses the pumping for $\tau_{w}=5.12$ ms. (b) Experimental data and numerical simulation of the LZ transition probability $\langle\it{P}_{T}\rangle$ (left axis) and the r.m.s. $S\text{-}T_{-}$ coupling strength $\langle\Delta_{ST}^{2}\rangle^{1/2}$ (right axis) as a function of LZ sweep number $N$. (c) Numerical results showing the synchronization of the nuclear spin precession around the external magnetic field $B^{z}$ between the two dots. The blue and red arrows denote the instantaneous orientation and relative amplitude of the in-plane components of $I_{1}$ and $I_{2}$ after the $N$th LZ sweep, where $\bm{I}_{i}$ is the average nuclear polarization in dot $i$. For demonstration purposes, nuclear spin dephasing and relaxation are not included in the simulation of (c).
Curiously, the $S\text{-}T_{-}$ transition probability appears to decrease with time. This is revealed by the decreasing signal amplitude at $\epsilon_{ST}$ over time [Fig. 1(c), right panel]. To explore this feature with better control, we perform DNP via Landau-Zener (LZ) sweeps through $\epsilon_{ST}$ [4, 5], [Fig. 1(b)]. We measure how the LZ transition probability depends on the number of sweeps. We repeat a “pump” cycle, consisting of a singlet initialization, an LZ sweep through $\epsilon_{ST}$, and a single-shot singlet-triplet measurement, 64 times [Fig. 2(a)]. During the LZ sweep, we sweep $\epsilon$ from 6.2 to 7.7 mV through $\epsilon_{ST}$ in $5$ $\mu$s, with a time interval $\tau_{w1}$ of $27~{}\mu$s to the next sweep. After this set of 64 pump cycles, we halt the pumping for 5.12 ms to allow the nuclear spin system to dephase and to implement auxiliary pulses to calibrate our measurement (see Supplemental Material). Figure 2(b) shows the triplet probability $\langle\it{P}_{T}\rangle$ after each LZ sweep as a function of the sweep number, averaged over more than $10^{4}$ pump-pause cycles.
The experimental data reveal that $\langle P_{T}\rangle$ rapidly decreases with the sweep number and saturates at a small value [Fig.2(b)]. The remainder of this paper will explore this phenomenon. One possible explanation for this effect is that our singlet-triplet readout contrast diminishes with increased nuclear polarization. Such an effect has been observed in past work in GaAs quantum dots [29, 30]. This effect occurs because the $T_{0}$ state relaxation rate increases significantly when the longitudinal magnetic gradient $\Delta E_{Z}$ becomes large, usually on the order of hundreds of MHz. However, this effect is unlikely to explain our data, because the typical $\Delta E_{Z}$ values are on the order of MHz in our device. In addition, we measure the $T_{-}$ state, which does not experience increased relaxation at large $\Delta E_{Z}$. Another possible explanation for our data is that the process of DNP is so efficient that $\epsilon_{ST}$ moves outside of our LZ sweep window. As discussed further below, and as shown in Fig. 4, this is also unlikely.
A third explanation is the formation of the nuclear-spin dark state. The dark state should emerge during DNP as the nuclei are driven into a state that decouples from the electrons [10, 14]. In the case of a double quantum dot, the dark state is characterized by a vanishing net transverse hyperfine gradient ($2\Delta_{ST}$ in magnitude). Figure 2(b) shows the root-mean-square values of $\Delta_{ST}$ extracted from experimental data, using the analytical expression for the LZ transition probability [31, 32] as $\langle P_{T}(t)\rangle=\langle 1-\exp[-\frac{(2\pi)^{2}\Delta_{ST}^{2}(t)}{\beta}]\rangle\approx\frac{(2\pi)^{2}\langle\Delta_{ST}^{2}(t)\rangle}{\beta}$, where $\beta$ is the sweep rate [Table 1]. $\langle\Delta_{ST}^{2}\rangle^{1/2}$ rapidly decreases with the LZ sweep number from 110 to 30 KHz.
The dark state can be understood in the following semiclassical picture. In thermal equilibrium, the average polarization along any direction of an ensemble of $N$ nuclear spins is approximately $\sqrt{N}$. Thus, the average imbalance in the transverse polarization between two dots with separate nuclear spin ensembles is also of order $\sqrt{N}$. During an LZ sweep, the quantum spin state of the electrons creates an effective magnetic field (the Knight field) that causes the difference of the transverse nuclear polarizations to rotate out of the $x$-$y$ plane and toward the $z$ direction. Thus, after about $\sqrt{N}$ LZ sweeps, the difference in the transverse polarization should vanish, leading to a scenario where $\Delta_{ST}=0$. In this “dark state”, the transverse polarizations of the two dots are the same, and they undergo a synchronized precession around the external field.
We perform numerical simulations of the full electron-nuclear system using the semiclassical techniques described in Refs. [10, 12] and the parameters in Table 1. This approach combines a quantum description of the two electron spins with a semiclassical model of the coherent dynamics for each of the nuclear spins interacting with the electrons in the quantum dots. At each time during the LZ sweeps, we compute the quantum state of the electrons, the effective Knight field exerted by the electrons on the nuclear spins, and how the nuclei evolve in this Knight field. We account for nuclear spin dephasing and relaxation. A nuclear spin relaxation time of $T_{1,n}=5$ s is chosen to reproduce the saturation $\langle P_{T}\rangle$ value in Fig. 2(b), and is broadly consistent with other measurements we have made. We also account for electrical noise experienced by the electrons. We consider 8000 and 4000 nuclei for dot 1 and dot 2, respectively, estimated according to the physical properties of the Si quantum dots. We use an hyperfine interaction strength of $|\bar{A}|=0.052$ GHz, corresponding to a two-electron spin dephasing time of approximately 0.8 $\mu$s [33]. The negative prefactor of $\bar{A}$ is determined by the signs of the electron and nuclear $g$-factors in Si. See the Supplemental Material for further details on our simulations.
Table 1: Parameters used for the simulations, where $N_{1(2)}$ is the number of nuclear spins in dot 1(2), $\bar{A}$ is the hyperfine interaction strength, $I$ is the amplitude of each nuclear spin, $\it{T}^{*}_{2,n}$ is the nuclear spin dephasing time, $\it{T}_{1,n}$ is the nuclear spin relaxation time, $\sigma_{\epsilon}$ describes the effect of electrical noise on $\epsilon$, and $\beta$ is the LZ sweep rate. We consider white noise for the nuclear spin dephasing unless otherwise indicated.
$N_{1}$ 8000 $\it{T}^{*}_{2,n}$ 4.1 ms
$N_{2}$ 4000 $\it{T}_{1,n}$ 5 s
$\bar{A}$0.052 GHz $\sigma_{\epsilon}$ $4.6\times 10^{-4}$ GHz
$I$ $1/2$ $\beta$ $2.8\times 10^{-6}$ GHz/ns
Figure 2(c) shows a simulated evolution of the nuclear spins during repeated LZ sweeps starting from a random state, confirming the semiclassical picture for the dark state formation. Figure 2(b) shows simulations of our experiment, which agree quantitatively with our measurements. The residual $\Delta_{ST}$ is associated with the finite time interval $\tau_{w1}$ between LZ sweeps (see Supplemental Material). One surprising feature of these data, which is captured by the simulations, is that $\Delta_{ST}$ reduces over about 10 sweeps, while the expected number of sweeps to reach the dark state is about $\sqrt{10^{4}}=10^{2}$, as shown in [Fig. 2(c)]. The explanation for this effect is that over the roughly $10^{4}$ pump-pause cycles (each of which contains 64 LZ sweeps), a significant longitudinal nuclear polarization builds up, effectively reducing the transverse polarization that reappears during the pause periods. The reduction of the transition probability in a number of sweeps reflecting the difference in transverse polarizations between the dots is a key feature of the nuclear-spin dark state.
We next investigate what happens during the pause periods. The semiclassical picture of the dark state provided above suggests that the synchronized precession of the net polarizations in each dot underlies vanishing electron-nuclear coupling. In this case, the primary mechanism for the nuclear spin system to exit the dark state involves inhomogeneous nuclear spin dephasing. The key idea is that after the characteristic dephasing time, a new, random transverse polarization of magnitude $\sqrt{N}$ appears in each dot. One might intuitively expect that nuclear spin relaxation should spoil the dark state. While the nuclear spin $T_{1}$ time does limit the lifetime of the dark state, dephasing occurs more rapidly.
Figure 3: Dark state dephasing. (a) Pulse sequence for the measurement consists of 256 pump cycles, a pause for a variable time of $t_{n}$, a probe cycle to measure $P_{T}$, and an additional pause that maintains the total duration of the sequence. (b) Measured transition probability $\langle P_{T}\rangle$ as a function of $t_{n}$ (filled circles). The data are fitted to the form $P_{0}(1-\exp[-(t_{n}/T_{c})])$ for nuclear spin dephasing due to white noise, with $P_{0}$ and $T_{c}$ as fitting parameters (solid line). The fitted value of $T_{2}^{*}=2T_{c}$ with the standard deviation is indicated. A numerical simulation considering white noise with the fitted $T_{2}^{*}$ time is also shown (dashed line). (c) The same data are fitted to the form $P_{0}(1-\exp[-(t_{n}/T_{c})^{2}])$ for dephasing due to quasi-static noise (solid line). The fitted value of $T_{2}^{*}=\sqrt{2}T_{c}$ with the standard deviation is indicated. The corresponding numerical simulation is shown (dashed line).
We explore this effect using a “pump-pause-probe-pause” sequence, as shown in Fig. 3(a). The sequence starts with 256 pump cycles to drive the nuclear spins into the dark state. Then, we let the nuclei precess and dephase freely for a variable time of $t_{n}$ without pumping. Immediately afterward, we perform one more LZ sweep to measure $\langle P_{T}\rangle$. A second “pause” is applied at the end to compensate for the variation in $t_{n}$, to keep the total duration of the pulse sequence the same for the purpose of consistency. Each probability value is obtained from over $10^{3}$ repetitions of such a pulse sequence. As shown in Fig. 3b, the measured $\langle P_{T}\rangle$ increases with increasing $t_{n}$ (filled circles), revealing a recovery of the $S\text{-}T_{-}$ coupling over the timescale of a few milliseconds, consistent with previous measurements of dephasing times of 29Si nuclei in Si quantum-dot devices [34].
To model these data, we consider the evolution of $\Delta_{ST}$ under the influence of nuclear spin dephasing. Assuming that at $t_{n}=0$, $\Delta_{ST}=0$, we analytically find that $\langle\Delta_{ST}^{2}\rangle\propto 1-\exp[-(t_{n}/T_{c})^{\gamma}]$. The characteristic time is closely related to the nuclear spin dephasing time, given by $T_{c}=T^{*}_{2,n}/\alpha$, with a factor $\alpha$ difference. Assuming that each nuclear spin experiences independent, Gaussian magnetic noise with a white power spectrum, $\langle\Delta_{ST}^{2}\rangle$ follows an exponential function with $\gamma=1$ and $\alpha=2$. For quasi-static noise, $\langle\Delta_{ST}^{2}\rangle$ follows a Gaussian function with $\gamma=2$ and $\alpha=\sqrt{2}$. By incorporating the form of $\langle\Delta_{ST}^{2}\rangle$ into the expression for the LZ transition probability $\langle P_{T}(t)\rangle\approx\frac{(2\pi)^{2}\langle\Delta_{ST}^{2}(t)\rangle}{\beta}$, we fit the experimental data. In the case of white noise [Fig. 3(b)], the best-fit value of $T^{*}_{2,n}$ is 4.1 ms, while for quasistatic noise [Fig. 3(c)], the best-fit value is 2.4 ms. In the figures, we also show numerical simulations with the fitted $T^{*}_{2,n}$ values. White noise appears to better describe the experimental data [Fig. 3(b)].
Figure 4: Increasing the electronic lifetime. (a) Pulse sequence for measurement with and without pumping. In the case without pumping, the LZ sweep in the pump cycles is replaced by a wait pulse at $\epsilon=0$. For the probe cycles, we initialize the singlet state and pulse to different values of $\epsilon$ for a variable evolution time $t_{e}$. (b) Measured $\langle P_{T}\rangle$ for $\epsilon$ values near $\epsilon_{ST}$ in the absence of pumping. (c) Measured $\langle P_{T}\rangle$ for $\epsilon$ values near $\epsilon_{ST}$ with pumping. (d) $\langle P_{T}\rangle$ at $\epsilon_{ST}$ as a function of $t_{e}$. The experimental data (filled triangles and circles) are extracted from (b) and (c), respectively, corresponding to the $\langle P_{T}\rangle$ maximum for each $t_{e}$ value. Numerical simulations for both cases are presented.
One major potential benefit of the nuclear-spin dark state is the possibility of increased electron-spin lifetimes, due to the reduced hyperfine coupling. To test this, we measure the relaxation time of the singlet state near $\epsilon_{ST}$ with and without DNP. In the former case, the pulse sequence starts with 64 pump cycles to form the dark state, followed by a “probe” cycle, as shown in Fig. 4(a). In the case without pumping, we use an identical pulse sequence but replace each LZ sweep in the pump cycle by a wait at $\epsilon=0$ mV, away from $\epsilon_{ST}$. In each probe cycle, with the electronic state initialized as a singlet, we pulse $\epsilon$ abruptly to a value near $\epsilon_{ST}$, wait at that $\epsilon$ for a variable time $t_{e}$, allowing the state to relax, and then abruptly pulse $\epsilon$ back to 0 mV for the measurement. Again, each probability value is obtained from over $10^{3}$ repeated measurements.
Figures 4(b) and (c) show the measured $\langle P_{T}\rangle$ after the wait near $\epsilon_{ST}$ for variable time $t_{e}$ without and with pumping, respectively. Without pumping, $\langle P_{T}\rangle$ at $\epsilon_{ST}$ increases over a time-scale of about a microsecond and saturates between 0.1 and 0.2, as indicated by the yellow region in Fig. 4(b) and filled triangles in Fig. 4(d). The relatively low saturation value reflects both the relatively large voltage resolution used in this measurement, as well as the presence of detuning noise in our system.
With pumping, however, $\langle P_{T}\rangle$ remains nearly zero up to at least 10 $\mu$s [Fig. 4(d)], indicating strongly suppressed electronic-spin relaxation. In principle, $\langle P_{T}\rangle$ should remain small for up to the lifetime of the dark state. Future experiments will focus on verifying this lifetime enhancement. As above, our numerical simulations agree well with our data [Fig. 4(d)]. We also note that even in the dark state, Fig. 4(c) shows that $\epsilon_{ST}$ still falls within our LZ sweep window, as discussed above.
IV Discussion
We have presented evidence for a nuclear-spin dark state in silicon, including the rapid reduction in electron-nuclear coupling strength with repeated Landau-Zener sweeps, measurements of the dark-state lifetime that are consistent with nuclear-spin dephasing, and measurements of significantly increased electron-spin lifetimes in the dark state. Together with these measurements, we also developed semiclassical numerical simulations that compute the behavior of all electron and nuclear spins involved in our device.
While our numerical predictions agree well with our data, further experimental work is required to determine the nuclear spin relaxation time $T_{1,n}$. Experimentally, we have observed that this relaxation time seems to depend on the exact configuration of the nuclear spins, as well as the pulses used to manipulate the electrons. Based on our simulations, it appears that the level of polarization achieved in the nuclear spin system depends sensitively on the relaxation rate.
Future work will focus on understanding the role of spin-orbit coupling in this system, and its relationship to a predicted topological phase transition [15]. Additional experiments will focus on how this understanding of the dark state can enable more efficient DNP, and possibly improved coherence of electronic spin-states. The extensive simulations we have presented may also shed light on the long-standing question of the origin of magnetic-gradient pumping in gate-defined quantum dots. More broadly, future research will seek to understand quantum effects in the nuclear spin system, and how these effects impact the electron spins. In addition, we plan to explore how to use the dark state as a quantum memory, resource for sensing, or other quantum applications [16, 17].
Note added: During the completion of this work, we became aware of related work exploring nuclear-spin dark states as quantum memories [35].
V Acknowledgments
We thank Lisa F. Edge of HRL Laboratories, LLC. for the epitaxial growth of the SiGe material and Elliot J. Connors for device fabrication. This work was sponsored by the Army Research Office under grants W911NF-17-1-0260 and W911NF-19-1-0167, and by the National Science Foundation under grant OMA 1936250 and DMR-1941673. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.